OpenStudy (liizzyliizz):

differential equations by separation of variables dy/dx= 4y (help with how to solve the differential equations)

OpenStudy (liizzyliizz):

i know how to do them, but i dont know how to check on the calculator and im not sure if i am doing them correctly.

OpenStudy (2bornot2b):

\[\int\limits_{}^{}(\frac{1}{4y})dy=\int\limits_{}^{}dx\]

OpenStudy (mathmate):

To check, substitute the function you got, differentiate with respect to x and see if you get 4y.

OpenStudy (jamesj):

if you want to verify your result, you don't use a calculator. You use pencil and paper. The very best way to verify you have the right result to a differential equation is to substitute your solution into the differential equation and show that it satisfies it.

OpenStudy (wasiqss):

james check out my prob too http://openstudy.com/study#/updates/4f0d9f5de4b084a815fcdc36

OpenStudy (jamesj):

btw, what solution did you find to this equation?

OpenStudy (anonymous):

dy/dx=4y 1/4y dy=dx ∫▒〖1/4y dy=∫▒1 dx〗 1/4 lny=x+c lny^(1/4)=x+c

OpenStudy (anonymous):

\[ \frac {dy}{dx}=4y\] \[\frac {dy}{4y}=dx\] \[\int\limits \frac {1}{4y}dy=\int\limits dx\] \[\frac {1}{4}\ln y=x+C\] \[\ln y=4x+C _{1}\] \[y=e ^{4x+4C _{1}}\] \[y=e ^{4x}e ^{4C _{1}}=e ^{4x}C\]

OpenStudy (anonymous):

4 of 1/4 cannot be taken to the RHS it will become the power of ln's variable

OpenStudy (anonymous):

I didn`t get what your saying

OpenStudy (anonymous):

1/4lny becomes lny^(1/4)

OpenStudy (jamesj):

Cinar's solution is correct.

OpenStudy (anonymous):

but as far as i have studied the co-effecient of lny becomes the power of y.can u explain?

OpenStudy (jamesj):

Yes, that's right. \[ a \ln b = \ln b^a \] But because we want to solve for y, we don't take that step, or if we do, it's just an intermediate of one, until we get to y is a function of x. That is what cinar has done.

OpenStudy (anonymous):

okay.got it

OpenStudy (anonymous):

tmrk, you can go on to solve problem like this, \[\ln y ^{1/4}=x+c => y ^{1/4}=e ^{x+c} => ( y ^{1/4})^{4}=(e ^{x+c})^{4}\] \[y=e ^{4x+4c} => y=e ^{4x}e ^{4c} =>y=e ^{4x}C\]

OpenStudy (anonymous):

thanks i got it.